# WorldCuisines: A Massive-Scale Benchmark for Multilingual and Multicultural Visual Question Answering on Global Cuisines

### v2

@@ 1 Introduction
As a result, similar food concepts can be found across different countries, reflecting a shared human culinary experience. Researchers use food as a proxy to model and analyze cultural dynamics, helping to quantify cultural differences across regions Adilazuarda et al. (2024). Many cultures have developed their own versions of “stuffed pasta” or “dumplings”, each with unique ingredients and preparation methods, often known by different names Gallani (2015), as illustrated in Figure 1. Small details like how the dumpling is shaped can signal the cultural background. Conversely, some dishes share the same name but have different meanings; for example, “jelly” in the U.S. refers to a fruit spread, while in the U.K. and parts of Asia, it refers to a gelatinous dessert Poppe (1992); Abe (2013). This culinary diversity presents a challenge for Vision Language Models (VLMs), which must accurately recognize and differentiate food items based on cultural context for applications like food recognition. These models navigate the complexities of names, ingredients, and preparation styles that vary widely across regions. VLMs have shown effectiveness in text captioning Liu et al. (2024b, c) and have been adapted to support multiple languages Geigle et al. (2023); Shin et al. (2024).
@@
  • We present WorldCuisines, the first massive scale benchmark consisting of 1 million high-quality multilingual and multicultural text-image pairs annotated by native speakers in their local languages. We publicly release our resources, i.e. datasets, [^1] We release WC-VQA at https://huggingface.co/datasets/worldcuisines/vqa and WC-KB consisting food, location, cuisine, and prompt templates at https://huggingface.co/worldcuisines. code, [^2] We release our code at https://github.com/worldcuisines/worldcuisines. and leaderboard [^3] We release our leaderboard at https://huggingface.co/spaces/worldcuisines/worldcuisines. to support and advance future research in this rapidly evolving field.
@@ 2.2.2 Metadata Annotation
Table 3: Accuracy (%) results of WC-VQA for Test Large (60k). MCQ and OEQ indicate multiple-choice question and open-ended question, respectively. Best and second-best are bolded and underlined, respectively. We employ an optimized prompt provided by the authors (see Subsection D.1 in the Appendix for further details).
@@ 2.2.3 Quality Assurance
Before beginning the quality assurance process, we first identify common issues that arise during the annotation and develop automated rules to detect easily identifiable annotation errors, such as incorrect string formatting. Annotators are then asked to correct these errors. To further ensure data quality and validity, we conduct several rounds of quality assurance. Initially, we focus on image quality by removing instances where images are blurry, dark, or contain distracting elements such as people or other dishes. We also verify image licenses by cross-referencing them with information on Wikimedia Commons. Next, we refine dish categorization and descriptions, ensuring consistency in category assignments and maintaining descriptions free from “information breaches” (e.g., excluding regional details from the description). We standardize cuisine names and eliminate any redundancies. Finally, we meticulously review all country and area information to ensure its accuracy. This comprehensive approach guarantees the integrity and reliability of our dataset.
@@ 2.3.2 Question and Context Construction
Dish name prediction (Task 1) is divided into three question variations depending on the context: (1a) no-context question, where we simply ask for the name of the dish without any provided context; (1b) contextualized question where we provide additional information related to cuisine or location; and (1c) adversarial contextualized question which are similar to the contextualized questions but may include misleading location information to assess the model’s robustness to irrelevant details. Only basic question without any provided context is available for regional cuisine prediction (Task 2). The data statistics for each task are presented in Table 2.
@@
Figure 5: Accuracy (%) categorized by language (left), language vitality (center), and language family (right). We classify the language vitality by following the classification from Joshi et al. (2020).
@@ 2.3.3 Multiple Language Translation
All questions and contexts are initially collected in English, which are then carefully translated by native human speakers into 30 language varieties: 23 different languages with 7 languages having two different varieties each. We instructed the translators to prioritise the naturalness, and then followed by the diversity of translations when the duplication occurs.
@@ 2.3.4 Generating VQA Triplets
To ensure no overlap in train and test subsets, we split the dishes and the multilingual-questions into two subsets each, to ensure no dish or multilingual questions leakage between train and test. For every subset, we apply random sampling to get a pair of dish and its multilingual-questions. We use the dish entry in our KB dataset to pick the image and the location to be injected to the context, if any. The answer candidates for multiple-choice were picked by utilizing similarity search (Section 2.3.1). We repeat this process until we reach the desired number of training or test samples, or until all possible dish and question combinations are used, discarding any duplicates.
@@ 4.4 Scaling Law
It is evident that models with large sizes perform better than smaller ones, showing the scaling law still exists in this experiment, as shown in Figure 6. It is very interesting to see the same trend across different model families (e.g., Llava, Qwen, and even GPT-4o series). However, it is pretty clear for open-source models, Llama 3.2 Instruct has the lead for overall performance, which may be due the coverage of multilingual data used in its training, although it is still unclear since there is no evidence or supporting information that can back up the finding. Regardless, NVLM-D model does not perform as good as their base model Qwen2 VL Instruct in our benchmark. One reason could be the NVLM model is highly tuned for English, but not in languages other than English.
@@ 5 Related Work
Several prior studies have focused on developing culturally relevant VQA benchmarks, including FM-IQA Gao et al. (2015), MCVQA Gupta et al. (2020), xGQA Pfeiffer et al. (2022), MaXM Changpinyo et al. (2023), MTVQA Tang et al. (2024), MABL Kabra et al. (2023), MAPS Liu et al. (2024a), and MaRVL Liu et al. (2021). Additionally, CVQA Romero et al. (2024) and CulturalVQA Nayak et al. (2024) provide VQA datasets that cover various regions and diverse topics, including food, with CVQA also offering questions in multiple languages alongside English translations. SEA-VQA Urailertprasert et al. (2024) specifically benchmarks the South East Asian region. In contrast, FoodieQA Li et al. (2024b) and World Wide Dishes Magomere et al. (2024) are focused exclusively on food-related benchmarks. Our work is similarly motivated by using food as a cultural proxy, but it distinguishes itself with a significantly larger dataset and broader coverage of languages and dialects.
@@ 6 Conclusion
We introduce WorldCuisines, an open-source, large-scale benchmark designed for multilingual and multicultural, visually grounded language understanding. It comprises over 1 million data points across 30 languages and dialects. Our findings reveal that this benchmark remains challenging for models, particularly with dishes from specific regions and in low-resource languages. This provides insight into how well models understand regional cuisines. To enhance usability, we offer a dedicated evaluation split with two datasets of varying sizes. Our evaluation shows that while VLMs perform better with the correct context, they struggle with adversarial contexts intended to mislead them. Additionally, we are releasing a comprehensive knowledge base, VQA dataset, code, and leaderboard as open-source resources to support future research.
@@
Regarding the food entries coverage, it is important to note that our food entries are currently sourced from English Wikipedia. Although we aim to include as many diverse dishes as possible, we acknowledge that this approach limits the coverage of some regions. Nevertheless, our dataset serves as a valuable starting point. In future work, we plan to incorporate entries from non-English Wikipedia pages to improve regional representation and cultural diversity. For evaluation purposes, we include accuracy metrics for overall model performance and BERTScore for more detailed analysis. However, we recognize that evaluating VQA model performance on multicultural data remains an open challenge. Appropriate evaluation metrics are needed to effectively model the diversity of cultural contexts and linguistic variations. Addressing this issue will be a key focus of our future research efforts.
@@ Appendix B Open-Source Collaborative Effort
Appendix C More Results
C.1 Primary Metric: Accuracy (%)
@@
C.2 Secondary Metric: BERTScore
@@
Table 7: Accuracy (%) results of WC-VQA. MCQ and OEQ indicate multiple-choice question and open-ended question, respectively. Best and second-best are bolded and underlined, respectively. We employ an optimized prompt provided by the authors (see Subsection D.1 in the Appendix for further details).
@@
Appendix D Evaluation
D.1 Prompt Sensitivity

### v4

@@ 1 Introduction
As a result, similar food concepts can be found across different countries, reflecting a shared human culinary heritage. Researchers use food as a proxy to model and analyze cultural dynamics, helping to quantify cultural differences across regions Adilazuarda et al. (2024). Many cultures have developed their own versions of “stuffed pasta” or “dumplings”, each with unique ingredients and preparation methods, often known by different names Gallani (2015), as illustrated in Figure 1. Small details like how the dumpling is shaped can signal the cultural background. Conversely, some dishes share the same name but have different meanings; for example, “jelly” in the U.S. refers to a fruit spread, while in the U.K. and parts of Asia, it refers to a gelatinous dessert Poppe (1992); Abe (2013). This culinary diversity presents a challenge for Vision Language Models (VLMs), which must accurately recognize and differentiate food items based on cultural context for applications like food recognition. These models navigate the complexities of names, ingredients, and preparation styles that vary widely across regions. VLMs have shown effectiveness in text captioning Liu et al. (2024b, c) and have been adapted to support multiple languages Geigle et al. (2023); Shin et al. (2024).
@@
  • We present WorldCuisines, the first massive scale benchmark consisting of 1 million high-quality multilingual and multicultural text-image pairs annotated by native speakers in their local languages. We publicly release our resources, i.e., datasets,  [^1] We release WC-VQA at https://huggingface.co/datasets/worldcuisines/vqa and WC-KB consisting food, location, cuisine, and prompt templates at https://huggingface.co/worldcuisines. code, [^2] We release our code at https://github.com/worldcuisines/worldcuisines. and leaderboard [^3] We release our leaderboard at https://huggingface.co/spaces/worldcuisines/worldcuisines. to advance future research in this rapidly evolving field.
@@ 2.2.2 Metadata Annotation
Table 3: Accuracy (%) results of WC-VQA for Test Large (60k). MCQ and OEQ indicate multiple-choice question and open-ended question, respectively. Best and second-best are bolded and underlined, respectively. We employ an optimized prompt provided by the authors (see Subsection E.1 in the Appendix for further details).
@@ 2.2.3 Quality Assurance
Before starting the quality assurance process, we first identify common issues that arise during the annotation and develop automated rules to detect easily identifiable annotation errors, such as incorrect string formatting. Annotators are then asked to correct these errors. To further ensure data quality and validity, we conduct several rounds of quality assurance. Initially, we focus on image quality by removing instances where images are blurry, dark, or contain distracting elements such as people or other dishes. We also verify image licenses by cross-referencing them with information on Wikimedia Commons. Next, we refine the dish categorization and descriptions, ensuring consistency in category assignments and maintaining descriptions free from “information breaches” (e.g., excluding regional details from the description). We standardize cuisine names and eliminate any redundancies. Finally, we meticulously review all country and area information to ensure its accuracy. This comprehensive approach guarantees the integrity and reliability of our dataset.
@@ 2.3.2 Question and Context Construction
Dish name prediction (Task 1) is divided into three question variations depending on the context: (1a) no-context question, where we simply ask for the name of the dish without any provided context; (1b) contextualized question where we provide additional information related to cuisine or location; and (1c) adversarial contextualized question which are similar to the contextualized questions but may include misleading location information to assess the model’s robustness to irrelevant details.
For example, consider coxinha from Brazil, shown in Figure 2 (1b). A query with additional context here would be: ‘‘What is the common name for this dish in Brazil?" Here, the origin of coxinha, Brazil, serves as the context. In contrast, adversarial context involves providing misleading or irrelevant information in terms of location or type of cuisine to assess the model’s robustness to such distractions. For instance, in the case of eggs benedict shown in Figure 2 (1c), an adversarial context would be: “Yesterday I had a nice lunch at a Korean restaurant. I am about to have this dish now. What is this dish called?” In this scenario, the model should ignore the irrelevant detail (“nice lunch at a Korean restaurant”) and focus solely on the image and the question.
Only basic question without any provided context is available for regional cuisine prediction (Task 2). The data statistics for each task are presented in Table 2.
@@
Figure 5: Accuracy (%) categorized by language (left), language vitality (center), and language family (right). We classify the language vitality by following the classification proposed by Joshi et al. (2020).
@@ 2.3.3 Multiple Language Translation
All questions and contexts are initially collected in English, which are then carefully translated by native human speakers into 30 language varieties: 23 different languages with 7 languages having two different varieties each. We instructed the translators to prioritize the naturalness, and then followed by the diversity of translations when the duplication occurs.
@@ 2.3.4 Generating VQA Triplets
To ensure no overlap in train and test subsets, we split the dishes and the multilingual-questions into two subsets each, to ensure no dish or multilingual questions leakage between train and test. For every subset, we apply random sampling to get a pair of dish and its multilingual-questions. We use the dish entry in our WorldCuisines KB dataset to pick the image and the location to be injected to the context, if any. The answer candidates for multiple-choice were picked by utilizing similarity search (Section 2.3.1). We repeat this process until we reach the desired number of training or test samples, or until all possible dish and question combinations are used, discarding any duplicates.
@@ 4.4 Scaling Law
It is evident that large models perform better than smaller ones, showing the scaling law still exists in this experiment, as shown in Figure 6. It is very interesting to see the same trend across different model families (e.g., Llava, Qwen, and even GPT-4o series). However, it is pretty clear for open-source models, Llama 3.2 Instruct has the lead for overall performance, which may be due the coverage of multilingual data used in its training, although it is still unclear since there is no evidence or supporting information that can back up the finding. Regardless, NVLM-D model does not perform as good as their base model Qwen2 VL Instruct in our benchmark. One reason could be the NVLM model is highly tuned for English, but not in languages other than English.
@@ 5 Related Work
Several prior studies have focused on developing culturally relevant VQA benchmarks, including FM-IQA Gao et al. (2015), MCVQA Gupta et al. (2020), xGQA Pfeiffer et al. (2022), MaXM Changpinyo et al. (2023), MTVQA Tang et al. (2024), MABL Kabra et al. (2023), MAPS Liu et al. (2024a), and MaRVL Liu et al. (2021). Additionally, CVQA Romero et al. (2024) and CulturalVQA Nayak et al. (2024) provide VQA datasets that cover various regions and diverse topics, including food, with CVQA also offering questions in multiple languages alongside English translations. SEA-VQA Urailertprasert et al. (2024) specifically benchmarks the South East Asian region. In contrast, FoodieQA Li et al. (2024b) and World Wide Dishes Magomere et al. (2024) are benchmark focusing on food. Our work is similarly motivated by using food as a cultural proxy, but it distinguishes itself with a significantly larger dataset and broader coverage of languages and cultures.
@@ 6 Conclusion
We introduce WorldCuisines, an open-source, large-scale benchmark designed for multilingual and multicultural, visually grounded language understanding. It comprises over 1 million data points across 30 languages and dialects. Our findings reveal that this benchmark remains challenging for VLMs, particularly with dishes from specific regions and in low-resource languages. This provides insight into how well models understand regional cuisines. To enhance usability, we offer a dedicated evaluation split with two datasets of varying sizes. Our evaluation shows that while VLMs perform better with the correct context, they struggle with adversarial contexts intended to mislead them. Additionally, we are releasing a comprehensive knowledge base, VQA dataset, code, and leaderboard as open-source resources to support future research.
Acknowledgements
@@
It is important to note that our food entries are currently sourced from English Wikipedia. Although we aim to include as many diverse dishes as possible, we acknowledge that this approach limits the coverage of some regions. This is due to language affects commonsense and its specific knowledge (Sakai et al., 2024), which in turns suggesting insufficiency of sourcing only English Wikipedia. Nevertheless, our dataset serves as a valuable starting point. In future work, we plan to incorporate entries from non-English Wikipedia pages to improve regional representation and cultural diversity. For evaluation purposes, we include accuracy metrics for overall model performance and BERTScore for more detailed analysis. However, we recognize that evaluating VQA model performance on multicultural data remains an open challenge. Appropriate evaluation metrics are needed to effectively model the diversity of cultural contexts and linguistic variations. Addressing this issue will be a key focus of our future research efforts.
@@ Appendix B Open-Source Collaborative Effort
Appendix C Detailed Dataset Construction
C.1 Dataset Compilation
Our dataset, comprising 2,414 dishes and 6,084 images, was meticulously compiled and verified manually. Key metadata includes dish name, alias, coarse- and fine-grained categories, cuisines, regions, descriptions, images, and their licenses. The compilation process followed these steps:
  • We listed dish names for annotation.
  • Annotators filled metadata fields and selected up to 8 licensed images per dish, guided by instruction documentation with examples for consistent accuracy.
  • Post-annotation, annotators formed subgroups to verify specific metadata categories, ensuring detailed and consistent data across fields such as categories, cuisines, regions, descriptions, and images.
C.2 Negative Sampling
Recall that from our annotations, we have detailed metadata for all 2,414 dishes, including the dish name, coarse-grained categories, fine-grained categories, countries, and text descriptions. The negative answers were sampled using the following procedure:
  • (1)
    We used a multilingual model, specifically E5-LARGE Instruct, to compute the text embeddings. Each embedding was generated by concatenating the dish name with its corresponding text description.
  • (2)
    To identify negative samples, we computed the cosine similarity between the embeddings of the target dish and those of all other dishes in the dataset. The top-K most similar dishes were selected under three different conditions:
    • Same Fine-grained Category: Select top-K dishes from the same fine-grained category as the target dish.
    • Same Coarse-grained Category: Select top-K dishes from the same coarse-grained category but potentially different fine-grained categories.
    • No Category Restriction: Select top-K dishes from the entire dataset without any restriction on categories.
    Here, we used K=15, resulting in 45 candidate dishes in total.
  • (3)
    Each MCQ consists of five options: one correct answer and four negative answers. The negative answers were chosen as follows:
    • Two Difficult Options: The first two negative answers were selected from dishes in the same fine-grained category. These are intended to be more challenging for the model to distinguish.
    • One Medium Option: The third negative answer was selected from dishes in the same coarse-grained category.
    • One Easy Option: The fourth negative answer was selected from dishes without any category restriction, making it likely to be easier to identify as incorrect.
    This approach ensures a balanced difficulty among the negative options, with two difficult, one medium, and one easy negative answer.
  • (4)
    (Optional) Specifically for task 2, where the question involves identifying the correct location (country) of a dish, we followed a slightly modified approach:
    • From the previously retrieved 4 negative options, we identified the countries associated with each dish.
    • We then excluded the countries that are valid locations for the correct dish. The remaining countries were used to create the negative options for the location-based question.
Appendix D More Results
D.1 Primary Metric: Accuracy (%)
@@
D.2 Secondary Metric: BERTScore
@@
Table 7: Accuracy (%) results of WC-VQA. MCQ and OEQ indicate multiple-choice question and open-ended question, respectively. Best and second-best are bolded and underlined, respectively. We employ an optimized prompt provided by the authors (see Subsection E.1 in the Appendix for further details).
@@
Appendix E Evaluation
E.1 Prompt Sensitivity